# GraphRAG-R1: Graph Retrieval-Augmented Generation with Process-Constrained Reinforcement Learning

### v1

@@ Abstract.
Graph Retrieval-Augmented Generation (GraphRAG) has shown great effectiveness in enhancing the reasoning abilities of Large Language Models (LLMs) by leveraging graph structures for knowledge representation and modeling complex real-world relationships. However, existing GraphRAG methods still face significant bottlenecks when handling complex problems that require multi-hop reasoning, as their query and retrieval phases are largely based on pre-defined heuristics and do not fully utilize the reasoning potentials of LLMs. To address this problem, we propose GraphRAG-R1, an adaptive GraphRAG framework by training LLMs with process-constrained outcome-based reinforcement learning (RL) to enhance the multi-hop reasoning ability. Our method can decompose complex problems, autonomously invoke retrieval tools to acquire necessary information, and perform effective reasoning. Specifically, we utilize a modified version of Group Relative Policy Optimization (GRPO) that supports rollout-with-thinking capability to train the model. Next, we design two process-constrained reward functions. To handle the shallow retrieval problem, we design a Progressive Retrieval Attenuation (PRA) reward to encourage essential retrievals. Then, to handle the over-thinking problem, we design Cost-Aware F1 (CAF) reward to balance the model performance with computational costs. We further design a phase-dependent training strategy, containing three training stages corresponding to cold start and these two rewards. These stages empower GraphRAG with format following, behavior shaping, and smartness optimization abilities, respectively. Lastly, our method adopts a hybrid graph-textual retrieval to improve the reasoning capacity. Extensive experimental results demonstrate that GraphRAG-R1 significantly boosts LLM capabilities in solving complex reasoning problems compared to state-of-the-art GraphRAG methods on both in-domain and out-of-domain datasets. Furthermore, our framework can be flexibly integrated with various existing retrieval methods, consistently delivering performance improvements.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Large Language Models (LLMs) have attracted ever-growing attention in the past few years due to their outstanding ability in answering complex questions and generating human-quality text (Vaswani et al., 2017; WANG et al., 2024). Despite remarkable progress, LLMs still exhibit significant limitations when faced with complex tasks that require integrating vast and diverse knowledge domains or require deep reasoning and multi-step logical inference (Talmor et al., 2019; Petroni et al., 2021). As shown in Figure 1(a), LLMs will provide an incorrect response when the question is over-complex and requires knowledge beyond their boundary. Graph Retrieval-Augmented Generation (GraphRAG) (Xiang et al., 2025; Zhang et al., 2025a; Peng et al., 2024; Han et al., 2024; Zhu et al., 2024; Li et al., 2025), which utilizes knowledge graphs for knowledge representation, enables effective modeling of complex real-world relationships, such as causality, hierarchy, and dependency (Edge et al., 2024; Peng et al., 2024).
However, existing GraphRAG methods still face significant performance bottlenecks when handling complex problems that require multi-hop reasoning. The main bottleneck is that existing GraphRAG methods rely on pre-defined or heuristic rules during query processing and graph retrieval, which are too rigid and inflexible. Concretely, during query processing, the semantic or lexical similarity search (Han et al., 2024; Gao et al., 2023b; Fan et al., 2024) remains the most widely adopted. For semantic similarity search, queries with high complexity are averaged during vectorization, making it difficult to gather accurate information for each distinct semantic element. On the other hand, lexical similarity search captures only the superficial meaning of the query, easily losing implicit relationships and reasoning logic. Therefore, as shown in Figure 1(b), existing GraphRAG methods are unable to answer the query, since the retrieval results are incomplete or incorrect when queries are overly complex. Besides, during the graph retrieval phase, current methods employ a combination of graph traversal and subgraph retrieval for path reasoning, which are rather rigid (Sun et al., [n. d.]; Ma et al., 2024; Jimenez Gutierrez et al., 2024; Guo et al., 2024; Li et al., 2024a). For instance, ToG (Sun et al., [n. d.]; Ma et al., 2024), a representative GraphRAG method, starts from a node in the knowledge graph and traverses the surrounding nodes based on breadth-first search strategy with pruning. While this strategy might work for short paths, it struggles with long-path problems, which require capturing logically sound but distant or even disconnected relationships. This inflexibility often leads to broken reasoning chains and failed retrievals, making it unsuitable for multi-hop problems with complex reasoning logic. As a result, the existing GraphRAG methods fail to fully leverage the LLM’s deep reasoning potential.
@@
To address this issue, we propose GraphRAG-R1, an adaptive GraphRAG framework with deep reasoning capabilities through process-constrained outcome-based RL. As shown in Figure 1(c), our method can decompose complex multi-hop problems by reasoning and autonomous, on-demand invocation of retrieval tools. Specifically, we employ a modified version of Group Relative Policy Optimization (GRPO) that supports rollout-with-thinking mechanism (Song et al., 2025) to autonomously leverage external retrieval tools on demand. Next, we design two process-constrained reward functions, Progressive Retrieval Attenuation (PRA) and Cost-Aware F1 (CAF), aiming to inject prior knowledge into the LLM’s training process. In detail, PRA aims to dynamically adjust the reward for retrieval actions. Initially, it encourages necessary retrievals to gather key information, preventing the model from shallow retrieval. As the reasoning deepens, PRA progressively attenuates retrieval rewards, guiding the model to rely more on acquired information for internal reasoning and avoid unnecessary retrieval overhead. On the other hand, CAF evaluates the final answer correctness using the F1 score while also accounting for the computational resources consumed during the reasoning and retrieval process. It encourages the model to achieve correct results more efficiently to prevent over-thinking. These two rewards collectively balance between effectiveness and computational cost. We further design a three-stage phase-dependent training strategy. In the first stage, we focus on the answer format to alleviate cold start problems. In the second and third stage, we focus on improving the answer quality using the aforementioned two rewards. Lastly, our method employs a hybrid graph-textual retrieval instead of only retrieving textual information as in existing GraphRAGs. By leveraging graph structures to precisely capture entity relationships while preserving the rich semantic details of the original text, our method provides more comprehensive and integrated knowledge for complex reasoning.
@@
  • We propose GraphRAG-R1, a pioneering framework that utilizes GRPO that supports rollout-with-thinking to enhance the reasoning ability of GraphRAG. To the best of our knowledge, we are the first to explore RL-based LLM training in GraphRAG.
@@
Figure 1. An example of comparing using only LLMs, GraphRAG, and GraphRAG-R1 in answering complex problems: (a) The LLM (Qwen-2.5-7b) directly produces an output, but the answer is incorrect, (b) HippoRAG2 (Gutiérrez et al., 2025) enhances the LLM (Qwen-2.5-7b) by retrieving external knowledge, but fail to generate correct outputs for complex problem, (c) our method successfully decomposes the problem and autonomously invokes retrieval methods, producing the correct output.
@@ 2.1. Graph Retrieval-Augmented Generation
Graph Retrieval-Augmented Generation (GraphRAG) offers a novel paradigm to enhance the reasoning abilities of LLMs by explicitly modeling entity relationships through graph structure such as knowledge graphs (Peng et al., 2024; Han et al., 2024). GraphRAG has been widely adopted across multiple domains. For instance, GraphRAG based on knowledge graph (Pan et al., 2024) incorporates techniques like subgraph retrieval and Graph Neural Networks (GNNs) (He et al., 2024) to enhance the accuracy and explainability of LLMs in complex knowledge reasoning question answering. GraphRAG based on document graphs (Sonawane and Kulkarni, 2014) model semantic relationships between documents and internal document structures, effectively supporting long-document summarization and cross-document information integration (Sarthi et al., [n. d.]). Furthermore, GraphRAG for specialized domains are increasingly trending. Examples include leveraging social graphs (Jiang and Ferrara, 2023; Zeng et al., 2024) to analyze user relationship for precision recommendations and constructing medical or disease graphs (Yang et al., 2024b; Li et al., 2024b) that integrate multi-source heterogeneous data to enable personalized treatment plan generation.
However, existing GraphRAG methods still face significant performance bottlenecks when facing complex reasoning problem, primarily due to pre-defined or heuristic rules during query processing and graph retrieval that struggle with highly complex multi-hop problems. For example, ToG (Sun et al., [n. d.]; Ma et al., 2024) introduces reasoning steps using breadth-first search (BFS) with pruning approach, but it heavily relies on prompt engineering and external LLMs and thus lacks flexibility. G-Retriever (He et al., 2024) adopts K-Nearest Neighbors-based retrieval with limited efficacy in multi-hop scenarios. Moreover, existing approaches suffer from shallow retrieval depth, since they often depend on predefined fixed paths or hop counts. HippoRAG (Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025) optimizes knowledge integration through personalized PageRank but and only supports single-hop retrieval. KGP (Wang et al., 2024) and LightRAG (Guo et al., 2024) are typically confined to 2-3 hops, exhibiting significant performance degradation when handling problems that require deep-dependence mining. In contrast, our work improves the reasoning ability of GraphRAG for complex problems using process-constrained RL, aiming to effectively resolves the issues of rigid retrieval strategies and complex reasoning shortcomings in existing GraphRAG approaches.
@@ 2.2. RL-Enhanced Retrieval Scheme
Recently, reinforcement learning (RL) has become increasingly important for complex reasoning (Xie et al., 2025; Yang et al., 2024a; Ouyang et al., 2022). For example, DeepSeek-R1 (Guo et al., 2025) shows that large-scale RL can endow LLMs with deep thinking capabilities. Some pioneering work have also adopted RL to enhance RAG (Jin et al., 2025; Song et al., 2025). For example, R1-Searcher (Song et al., 2025) and Search-R1 (Jin et al., 2025) achieve co-training of retrieval and generation, effectively improving the model’s ability to solve complex problems. However, a systematic exploration of RL schemes within the GraphRAG domain remains open. Technically speaking, directly extending existing RL methods into GraphRAG faces several challenges, Firstly, the existing methods usually employ a single outcome-oriented reward function, e.g., based solely on the correctness of final answer. However, this design becomes too simple for GraphRAG, since complex problem may involve multiple different retrieval calls to different components of the external knowledge graph. More critically, existing methods have often encountered with reward hacking in RL training, potentially resulting in shallow retrieval and over-thinking in GraphRAG. To tackle these issues, we propose GraphRAG-R1 with tailored reward functions and training strategy to non-trivially explore the potential of RL in enhancing GraphRAG, to the best of our knowledge, for the first time.
@@ 3.1. Rollout Retrieval Enhanced GRPO
where $\pi_{\text{ref}}(\cdot)$ is the reference LLM, $\mathbb{D}_{\text{KL}}(\cdot)$ is the KL-divergence measure, and $\beta$ is a hyperparameter that controls the strength of the KL-divergence regularization. More specifically, our method builds on the GRPO algorithm (Guo et al., 2025), which uses the average reward of multiple sampled outputs as a baseline. Specifically, for each input question $q$, GRPO samples a group of $G$ responses ${o_{1},o_{2},...,o_{G}}$ using the old policy $\pi_{\theta_{\text{old}}}$. The policy model is then optimized by maximizing the following objective function:
@@
where the group size $G$ determines the number of responses generated per query, $\theta_{\text{old}}$ denotes the previous policy parameters, $\epsilon$ is the clipping range hyperparameter to constrain policy updates, and $A_{i}$ represents the advantage, computed based on the relative rewards of outputs within each group as follow:
@@
To better adapt GRPO to our GraphRAG scenario, we adopt two key modifications proposed by R1-Searcher (Song et al., 2025): Rollout-with-Thinking and retrieval-masked loss. The core objective is to enable the model to autonomously leverage external retrieval tools on demand during inference, i.e., accessing supplementary knowledge when needed to solve problems beyond its internal capacity, and receive rewards for generating correct answers.

### v2

@@ Abstract.
Graph Retrieval-Augmented Generation (GraphRAG) has shown great effectiveness in enhancing the reasoning abilities of Large Language Models (LLMs) by leveraging graph structures for knowledge representation and modeling complex real-world relationships. However, existing GraphRAG methods still face significant bottlenecks when handling complex problems that require multi-hop reasoning, as their query and retrieval phases are largely based on pre-defined heuristics and do not fully utilize the reasoning potentials of LLMs. To address this problem, we propose GraphRAG-R1, an adaptive GraphRAG framework by training LLMs with process-constrained outcome-based reinforcement learning (RL) to enhance the multi-hop reasoning ability. Our method can decompose complex problems, autonomously invoke retrieval tools to acquire necessary information, and perform effective reasoning. Specifically, we utilize a modified version of Group Relative Policy Optimization (GRPO) that supports rollout-with-thinking capability to train the model. Next, we design two process-constrained reward functions. To handle the shallow retrieval problem, we design a Progressive Retrieval Attenuation (PRA) reward to encourage essential retrievals. Then, to handle the over-thinking problem, we design a Cost-Aware F1 (CAF) reward to balance the model performance with computational costs. We further design a phase-dependent training strategy, containing three training stages corresponding to cold start and these two rewards. These stages empower GraphRAG with format following, behavior shaping, and smartness optimization abilities, respectively. Lastly, our method adopts a hybrid graph-textual retrieval to improve the reasoning capacity. Extensive experimental results demonstrate that GraphRAG-R1 significantly boosts LLM capabilities in solving complex reasoning problems compared to state-of-the-art GraphRAG methods on both in-domain and out-of-domain datasets. Furthermore, our framework can be flexibly integrated with various existing retrieval methods, consistently delivering performance improvements.
Reinforcement Learning, GraphRAG, Knowledge Graph
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792589isbn: 979-8-4007-2307-0/2026/04submissionid: rfp3241ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Machine learning
Resource Availability: The source code has been made publicly available at https://doi.org/10.5281/zenodo.18349146 and can also be accessed via the project repository: https://github.com/ycygit/GraphRAG-R1. The model weights are available at https://doi.org/10.57967/hf/7629 and https://huggingface.co/yuchuanyue/GraphRAG-R1.
@@ 1. Introduction
Large Language Models (LLMs) have attracted ever-growing attention in the past few years due to their outstanding ability in answering complex questions and generating human-quality text (Vaswani et al., 2017; WANG et al., 2024). Despite remarkable progress, LLMs still exhibit significant limitations when faced with complex tasks that require integrating vast and diverse knowledge domains or require deep reasoning and multi-step logical inference (Talmor et al., 2019; Petroni et al., 2021). As shown in Figure 1(a), LLMs tend to provide an incorrect response when the question requires knowledge beyond their scope. Graph Retrieval-Augmented Generation (GraphRAG) (Xiang et al., 2025; Zhang et al., 2025a; Peng et al., 2024; Han et al., 2024; Zhu et al., 2024; Li et al., 2025), which utilizes knowledge graphs for knowledge representation, enables effective modeling of complex real-world relationships, such as causality, hierarchy, and dependency (Edge et al., 2024; Peng et al., 2024).
However, existing GraphRAG methods still face significant performance bottlenecks when handling complex problems that require multi-hop reasoning. The main bottleneck is that existing methods rely on pre-defined or heuristic rules during query processing and retrieval, which are too rigid and inflexible. Concretely, during query processing, the semantic or lexical similarity search (Han et al., 2024; Gao et al., 2023b; Fan et al., 2024) remains the most widely adopted. For semantic similarity search, queries with high complexity are averaged out during vectorization, making it difficult to gather accurate information for each distinct semantic element. On the other hand, lexical similarity search captures only the superficial meaning of the query, easily losing implicit relationships and reasoning logic. Therefore, as shown in Figure 1(b), existing GraphRAG methods are unable to answer the query, since the retrieval results are incomplete or incorrect when queries are overly complex. Besides, during the graph retrieval phase, current methods employ a combination of graph traversal and subgraph retrieval for path reasoning, which are rather rigid (Sun et al., [n. d.]; Ma et al., 2024; Jimenez Gutierrez et al., 2024; Guo et al., 2024; Li et al., 2024a). For instance, ToG (Sun et al., [n. d.]; Ma et al., 2024), a representative GraphRAG method, starts from a node in the graph and traverses the surrounding nodes based on breadth-first search strategy with pruning. While this strategy might work for short paths, it struggles with long-path problems, which require capturing logically sound but distant or even disconnected relationships. This inflexibility often leads to broken reasoning chains and failed retrievals, making it unsuitable for multi-hop problems with complex reasoning logic. As a result, the existing GraphRAG methods fail to fully leverage the LLM’s deep reasoning potential.
@@
To address this issue, we propose GraphRAG-R1, an adaptive GraphRAG framework with deep reasoning capabilities through process-constrained outcome-based RL. As shown in Figure 1(c), our method can decompose complex multi-hop problems by reasoning and autonomous invocation of retrieval tools. Specifically, we employ a modified version of Group Relative Policy Optimization (GRPO) that supports rollout-with-thinking mechanism (Song et al., 2025) to leverage external retrieval tools on demand. Next, we design two process-constrained reward functions, Progressive Retrieval Attenuation (PRA) and Cost-Aware F1 (CAF), aiming to inject prior knowledge into the LLM’s training process. In detail, PRA aims to dynamically adjust the reward for retrieval actions. It encourages necessary retrievals , preventing the model from shallow retrieval. As the reasoning deepens, PRA progressively attenuates retrieval rewards, guiding the model to avoid unnecessary retrieval overhead. On the other hand, CAF evaluates the final answer correctness using the F1 score while also accounting for the computational resources consumed during the reasoning and retrieval process. It encourages the model to achieve correct results more efficiently to prevent over-thinking. These two rewards collectively balance effectiveness and computational cost. We further design a three-stage phase-dependent training strategy. In the first stage, we focus on the answer format to alleviate cold start problems. In the second and third stages, we focus on improving the answer quality using the aforementioned two rewards. Lastly, our method employs a hybrid graph-textual retrieval. By leveraging graph structures to precisely capture entity relationships while preserving the rich semantic details of the original text, our method provides more comprehensive and integrated knowledge for complex reasoning.
@@
  • We propose GraphRAG-R1, a pioneering framework that utilizes GRPO that supports rollout-with-thinking to enhance the reasoning ability of GraphRAG.
@@
Figure 1. An example of comparing using only LLMs, GraphRAG, and GraphRAG-R1 in answering complex problems: (a) The LLM (Qwen2.5-7B) directly produces an output, but the answer is incorrect, (b) HippoRAG2 (Gutiérrez et al., 2025) enhances the LLM (Qwen2.5-7B) by retrieving external knowledge, but fail to generate correct outputs for complex problem, (c) our method successfully decomposes the problem and autonomously invokes retrieval methods, producing the correct output.
This figure provides a comparative schematic illustrating the performance of three distinct approaches in handling the same complex reasoning problem, visually demonstrating the superiority of the proposed GraphRAG-R1 framework. The schematic is divided into three juxtaposed sub-figures: (a) Using only the base Large Language Model (LLM, Qwen2.5-7B): The model directly generates an answer, which is incorrect. This reveals the limitations of LLMs regarding knowledge scope and multi-step reasoning. (b) Using an existing Graph Retrieval-Augmented Generation method (HippoRAG2) to augment the same LLM: While this method introduces external knowledge retrieval, it still fails to produce the correct answer for this complex problem. This highlights the bottlenecks of traditional GraphRAG methods in flexible problem decomposition and multi-hop reasoning. (c) Applying the proposed GraphRAG-R1 method: The model successfully decomposes the complex problem into sub-steps, autonomously and on-demand invokes external retrieval tools to acquire necessary information, and finally integrates the reasoning to produce the correct answer. Through this comparison, the figure clearly validates the effectiveness of GraphRAG-R1 in enhancing the LLM’s capability to solve complex problems.
@@ 2.1. Graph Retrieval-Augmented Generation
Graph Retrieval-Augmented Generation (GraphRAG) offers a novel paradigm to enhance the reasoning abilities of LLMs by explicitly modeling entity relationships through graph structure such as knowledge graphs (Peng et al., 2024; Han et al., 2024). GraphRAG has been widely adopted across multiple domains. For instance, GraphRAG based on knowledge graph (Pan et al., 2024) incorporates techniques like subgraph retrieval and Graph Neural Networks (GNNs) (He et al., 2024) to enhance the accuracy and explainability of LLMs in complex knowledge reasoning question answering. GraphRAG based on document graphs (Sonawane and Kulkarni, 2014) models semantic relationships between documents and internal document structures, effectively supporting long-document summarization and cross-document information integration (Sarthi et al., [n. d.]). Furthermore, GraphRAG methods for specialized domains are increasingly trending. Examples include leveraging social graphs (Jiang and Ferrara, 2023; Zeng et al., 2024) to analyze user relationships for precision recommendations and constructing medical or disease graphs (Yang et al., 2024b; Li et al., 2024b) that integrate multi-source heterogeneous data to enable personalized treatment plan generation.
However, existing GraphRAG methods still face significant performance bottlenecks when facing complex reasoning problems. For example, ToG (Sun et al., [n. d.]; Ma et al., 2024) introduces reasoning steps using breadth-first search (BFS) with pruning approach, but it heavily relies on external LLMs and thus lacks flexibility. G-Retriever (He et al., 2024) adopts K-Nearest Neighbors-based retrieval with limited efficacy in multi-hop scenarios. Moreover, existing approaches suffer from shallow retrieval depth, since they often depend on predefined fixed paths or hop counts. HippoRAG (Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025) only supports single-hop retrieval. KGP (Wang et al., 2024) and LightRAG (Guo et al., 2024) are typically confined to 2-3 hops, exhibiting performance degradation when handling problems that require deep-dependence mining. In contrast, our work improves the reasoning ability of GraphRAG for complex problems using process-constrained RL, aiming to effectively resolves the issues of rigid retrieval strategies and complex reasoning shortcomings in existing GraphRAG approaches.
@@ 2.2. RL-Enhanced Retrieval Scheme
Recently, reinforcement learning (RL) has become increasingly important for complex reasoning (Xie et al., 2025; Yang et al., 2024a; Ouyang et al., 2022). For example, DeepSeek-R1 (Guo et al., 2025) shows that large-scale RL can endow LLMs with deep thinking capabilities. Some pioneering works have also adopted RL to enhance RAG (Jin et al., 2025; Song et al., 2025). For example, R1-Searcher (Song et al., 2025) and Search-R1 (Jin et al., 2025) achieve co-training of retrieval and generation, effectively improving the model’s ability to solve complex problems. However, a systematic exploration of RL schemes within the GraphRAG domain remains open. Technically speaking, directly extending existing RL methods into GraphRAG faces several challenges. Firstly, the existing methods usually employ a single outcome-oriented reward function, e.g., based solely on the correctness of final answer. However, this design becomes too simple for GraphRAG, since a complex problem may involve complex retrieval calls of the external knowledge graph. More critically, existing methods have often encountered reward hacking in RL training, potentially resulting in shallow retrieval and over-thinking in GraphRAG. To tackle these issues, we propose GraphRAG-R1 with tailored reward functions and training strategy to non-trivially explore the potential of RL in enhancing GraphRAG.
@@ 3. Method
This figure presents a comprehensive architectural overview of the proposed GraphRAG-R1 framework, illustrating its three core, interconnected components: (a) Rollout Retrieval Enhanced GRPO as the LLM Training Strategy: This component depicts the reinforcement learning (RL) training paradigm, specifically a modified Group Relative Policy Optimization (GRPO) algorithm. It visually represents how the LLM policy is trained via process-constrained outcome-based RL, enabling the model to autonomously invoke retrieval tools during its reasoning ”rollout” or thought process. (b) Process-Constrained Reward Designs: This section details the design of the reward functions that guide and constrain the model’s behavior during RL training. It highlights the two novel rewards and illustrates the phase-dependent training strategy that applies them sequentially (cold start, behavior shaping, smartness optimization) to effectively shape the model’s reasoning and retrieval patterns. (c) Hybrid Graph-Textual Retrieval Module: This part contrasts the proposed retrieval mechanism with traditional text-only retrieval. It visually demonstrates how combining structured graph triplets (capturing precise entity relationships) with unstructured text snippets (preserving rich semantic context) provides more comprehensive and integrated knowledge, thereby supplying higher-quality information for the LLM’s complex reasoning tasks.
@@ 3.1. Rollout Retrieval Enhanced GRPO
where $\pi_{\text{ref}}(\cdot)$ is the reference LLM, $\mathbb{D}_{\text{KL}}(\cdot)$ is the KL-divergence measure, and $\beta$ is a hyperparameter that controls the strength of the KL-divergence regularization. More specifically, our method builds on the GRPO algorithm (Guo et al., 2025), which uses the average reward of multiple sampled outputs as a baseline. During training, for each input query $q$, GRPO samples a group of $G$ responses $o_{1},o_{2},...,o_{G}$} using the old policy $\pi_{\theta_{\text{old}}}$. The policy model is then optimized by maximizing the following objective function:
@@
where the group size $G$ determines the number of responses generated per query, $\theta_{\text{old}}$ denotes the previous policy parameters, $\epsilon$ is the clipping range hyperparameter to constrain policy updates, and $A_{i}$ represents the advantage, computed based on the relative rewards of outputs within each group as follows:
@@
To better adapt GRPO to our GraphRAG scenario, we adopt two key modifications proposed by R1-Searcher (Song et al., 2025): Rollout-with-Thinking and Retrieval-Masked Loss. The core objective is to enable the model to autonomously leverage external retrieval tools on demand during inference and receive rewards for generating correct answers.